# FullStack Bench: Evaluating LLMs as Full Stack Coders

### v2

@@ 4 Experiments
OpenCoder-1.5B-Instruct 20.47 38.71 43.00 45.94 23.08 41.84 22.50 11.03 24.60 20.00 42.65 34.57 33.08
Qwen2.5-Coder-1.5B-Instruct 17.44 33.87 40.00 35.03 27.27 44.56 18.75 10.66 43.65 15.00 48.53 32.45 31.59
DeepSeek-Coder-1.3B-Instruct 13.49 31.38 30.50 38.07 22.03 36.91 16.25 5.88 20.63 28.33 39.71 30.32 27.24
@@
Yi-Coder-9B-Chat 36.28 47.07 29.50 61.17 47.20 43.19 43.75 38.60 44.44 33.33 52.94 47.88 44.69
OpenCoder-8B-Instruct 33.02 48.24 30.00 58.12 32.52 42.34 36.25 26.47 46.03 36.67 44.12 38.83 41.08
DeepSeek-Coder-7B-Instruct-v1.5 36.28 45.90 27.00 56.09 36.01 41.33 27.50 19.85 46.83 41.67 39.71 42.55 40.22
DeepSeek-Coder-6.7B-Instruct 33.95 46.77 26.00 58.38 39.51 37.75 33.75 19.12 46.03 31.67 55.88 42.02 40.16
CodeQwen1.5-7B-Chat 36.51 43.25 36.00 48.73 25.17 40.64 23.75 16.18 43.65 41.67 41.18 40.43 37.76
CodeLlama-7B-Instruct 18.84 21.12 29.50 29.70 18.53 40.65 8.75 10.29 30.16 10.00 45.59 28.73 25.40
@@
Qwen2.5-Coder-14B-Instruct 48.60 59.09 42.50 64.97 65.03 51.53 47.50 26.47 53.97 51.67 66.18 56.38 53.41
DeepSeekCoder-v2-Lite-Instruct 46.05 57.33 41.50 56.35 53.85 44.22 38.75 20.59 49.21 38.33 47.06 51.59 47.69
StarCoder2-15B-Instruct-v0.1 33.95 42.08 37.00 61.42 35.31 46.77 41.25 26.84 53.17 30.00 44.12 50.00 42.68
CodeLlama-13B-Instruct 22.79 23.31 28.50 29.95 17.13 41.84 13.75 11.76 33.33 10.00 45.59 32.98 27.00
@@
DeepSeekCoder-v2-Instruct 52.09 64.22 44.00 69.80 70.98 46.09 50.00 38.60 53.97 51.67 64.71 61.17 56.37
Qwen2.5-Coder-32B-Instruct 34.65 60.70 44.00 68.02 69.58 52.38 56.25 19.49 57.94 51.67 60.29 52.66 52.40
DeepSeekCoder-33B-Instruct 40.23 51.32 35.00 64.72 50.35 46.26 47.50 39.34 54.76 46.67 50.00 45.75 48.19
CodeLlama-34B-Instruct 20.23 23.61 27.50 38.07 14.34 41.15 17.50 16.54 33.33 13.33 52.94 26.06 27.56
@@
Qwen2.5-72B-Instruct 51.16 61.73 37.00 62.18 72.73 45.92 46.25 19.85 56.35 41.67 58.82 51.06 52.19
Llama3.1-70B-Instruct 42.33 53.66 31.50 67.01 63.64 46.94 52.50 37.13 57.14 41.67 58.82 51.06 50.65
@@
OpenAI o1-preview 71.63 71.26 48.00 74.62 80.42 54.42 55.00 61.40 59.52 61.67 69.12 67.56 66.12
OpenAI o1-mini 68.60 74.78 43.50 68.53 79.72 46.59 52.50 36.03 54.76 58.33 64.71 64.36 61.44
Claude-35-Sonnet 59.30 66.87 43.00 71.57 75.87 45.75 57.50 61.76 54.76 61.67 67.65 62.23 60.70
GPT 4o-0806 59.53 68.18 40.00 72.34 77.62 47.45 67.50 31.25 57.14 51.67 64.71 60.63 58.89
Doubao-Coder-Preview 56.05 64.52 42.50 71.83 75.17 51.36 42.50 35.29 54.76 55.00 52.94 56.91 57.53
DeepSeek-v2.5 49.30 62.47 38.50 70.81 73.43 47.45 47.50 54.78 52.38 60.00 60.29 62.23 57.20
GLM-4-Plus 45.12 60.11 46.50 71.07 69.23 48.64 51.25 30.15 60.32 56.67 63.24 54.25 54.51
Qwen-Max 51.16 59.68 42.00 56.09 66.43 43.54 42.50 19.12 51.59 25.00 48.53 49.47 49.50
@@ 4.1 Experimental Setup
We chat with the close-sourced models via API calls in a similar manner, where top_k is set to 1 and max_new_tokens is set to 2048. The prompt template is kept the same and the input is not tokenized at our end.
After the model inference, the first code block formatted in Markdown is extracted from the generated output. If no Markdown code block is detected, a heuristic approach is employed to identify and extract incomplete code snippets. The extracted code, combined with predefined test cases, is then used to synthesize the complete code, which is subsequently evaluated for correctness.
@@ 4.2 Results and Analysis
As illustrated in Figure 5, we visualize the model performance on different domains in FullStack Bench. From Figure 5, we could find that the performance of AI coders varies significantly in BP, AP, MA, ML, and MM. The largest range occurs in MA, with the best mathematician being OpenAI o1-preview (80.42) while the worst is CodeLlama-34B-Instruct (14.34). Mathematical programming requires models to be proficient in both math and code, and those trained on a code-highly-concentrated corpus would struggle to achieve high scores in MA. Similarly, the variances of ML, SC, and MM, are also remarkable, as each of these problems requires domain knowledge beyond coding.
In Figure 5, we visualize the first place in each model division. The performance trends of Qwen2.5-Coder-7B-Instruct, Qwen2.5-Coder-14B-Instruct, DeepSeek-Coder-v2-Instruct, and Qwen2.5-72B-Instruct are generally consistent. Except for OpenAI o1-preview, the results of all models on SC lie in the trough. Besides, the curve of OpenCoder-1.5B-Instruct goes in a different direction than others, which indicates that the model might be trained on distinctive distributed data.
@@
OpenCoder-1.5B-Instruct 76.67 21.96 49.49 0.00 26.47 50.62 49.56 60.26 61.11 27.96 8.75 36.84 6.67 28.57 25.00 52.63 33.08
Qwen2.5-Coder-1.5B-Instruct 83.33 11.68 17.68 2.17 38.24 57.50 52.84 55.13 55.56 27.96 11.25 26.32 16.67 7.14 40.00 50.00 31.59
DeepSeek-Coder-1.3B-Instruct 53.33 14.02 43.43 2.17 26.47 43.12 35.59 58.97 63.89 24.17 5.00 23.68 10.00 5.36 22.50 47.37 27.24
@@
Yi-Coder-9B-Chat 76.67 24.77 74.24 10.87 32.35 68.12 27.73 73.08 69.44 45.62 26.25 55.26 50.00 42.86 55.00 65.79 44.69
OpenCoder-8B-Instruct 66.67 32.71 49.49 11.96 33.82 68.12 32.75 71.79 83.33 40.46 15.00 50.00 10.00 41.07 62.50 68.42 41.08
DeepSeek-Coder-7B-Instruct-v1.5 46.67 23.83 68.18 27.17 47.06 62.50 27.95 71.79 63.89 38.27 18.75 52.63 50.00 28.57 58.75 50.00 40.22
DeepSeek-Coder-6.7B-Instruct 73.33 28.97 61.11 5.43 35.29 65.00 20.96 67.95 86.11 40.64 28.75 31.58 38.33 35.71 58.75 68.42 40.16
CodeQwen1.5-7B-Chat 70.00 27.57 40.91 14.13 32.35 57.50 34.72 58.97 61.11 35.49 11.25 57.89 58.33 51.79 53.75 57.89 37.76
CodeLlama-7B-Instruct 70.00 13.08 46.46 11.96 13.24 51.88 35.81 37.18 41.67 19.14 10.00 15.79 23.33 16.07 33.75 47.37 25.40
@@
Qwen2.5-Coder-14B-Instruct 93.33 40.19 63.13 25.00 51.47 65.62 47.38 82.05 91.67 53.14 31.25 63.16 66.67 42.86 63.75 65.79 53.41
DeepSeekCoder-v2-Lite-Instruct 56.67 36.92 54.04 22.83 42.65 63.75 44.98 71.79 75.00 46.50 38.75 57.89 63.33 44.64 51.25 60.53 47.69
StarCoder2-15B-Instruct-v0.1 56.67 24.30 61.11 23.91 36.76 54.37 43.89 67.95 44.44 41.65 25.00 63.16 26.67 23.21 60.00 57.89 42.68
CodeLlama-13B-Instruct 66.67 11.21 54.55 10.87 26.47 55.00 33.41 43.59 38.89 20.32 13.75 44.74 21.67 25.00 33.75 44.74 27.00
@@
DeepSeekCoder-v2-Instruct 83.33 45.79 72.22 34.78 45.59 59.38 47.16 76.92 69.44 57.17 35.00 68.42 70.00 67.86 67.50 63.16 56.37
Qwen2.5-Coder-32B-Instruct 86.67 42.06 73.74 26.09 39.71 72.50 43.23 82.05 63.89 54.68 6.25 50.00 10.00 42.86 68.75 57.89 52.40
DeepSeekCoder-33B-Instruct 66.67 28.50 67.68 18.48 47.06 65.00 36.46 70.51 58.33 49.29 37.50 52.63 48.33 51.79 65.00 60.53 48.19
CodeLlama-34B-Instruct 83.33 12.62 42.93 10.87 27.94 60.00 31.66 42.31 61.11 22.27 7.50 31.58 21.67 14.29 43.75 47.37 27.56
@@
Qwen2.5-72B-Instruct 80.00 31.31 49.49 30.43 52.94 74.38 39.74 76.92 88.89 53.61 43.75 52.63 73.33 55.36 72.50 57.89 52.19
Llama3.1-70B-Instruct 80.00 33.18 71.21 38.04 38.24 78.75 29.26 75.64 38.89 52.67 33.75 73.68 61.67 32.14 70.00 63.16 50.65
@@
OpenAI o1-preview 90.00 52.80 81.31 50.00 45.59 74.38 40.61 78.21 69.44 70.14 70.00 92.11 90.00 89.29 71.25 68.42 66.12
OpenAI o1-mini 83.33 65.42 74.24 46.74 52.94 72.50 36.90 84.62 72.22 61.73 61.25 81.58 90.00 78.57 73.75 68.42 61.44
Claude-35-Sonnet 90.00 42.99 77.78 69.57 27.94 78.12 32.75 73.08 47.22 65.88 37.50 86.84 85.00 73.21 70.00 52.63 60.70
GPT 4o-0806 96.67 45.79 72.22 42.39 51.47 71.25 40.61 85.90 58.33 59.72 63.75 78.95 81.67 60.71 76.25 57.89 58.89
Doubao-Coder-Preview 66.67 40.19 73.74 33.70 38.24 68.13 37.12 83.33 75.00 60.96 53.75 71.05 71.76 75.00 65.00 65.79 57.53
DeepSeek-v2.5 80.00 45.33 75.25 23.91 42.65 73.12 39.30 80.77 80.56 59.77 41.25 63.16 66.67 60.71 67.50 68.42 57.20
GLM-4-Plus 83.33 35.98 73.74 26.09 30.88 69.38 52.62 85.90 91.67 53.91 22.50 50.00 63.33 51.79 73.75 55.26 54.51
Qwen-Max 66.67 34.11 35.86 32.61 41.18 74.38 41.27 76.92 72.22 50.18 40.00 63.16 71.67 51.79 70.00 60.53 49.50
@@ 4.3 Analysis on the performance of different programming languages
We present results of different programming languages in Table 3 and Figure 6. The majority of the AI coders are able to solve Bash programming well. Whereas the performance of C++, C#, and Ruby varies more, suggesting that they are selectively included in the training corpus by the model designers. Worser, some coders, especially those 1B+ tiny models behave poorly on D, R, and Scala as their pass rates are lower than 10%.
Our SandboxFusion provides feedback from the compilers. We evaluate the compilation pass rates of model responses in compiled languages including C++, D, Java, Rust, and Scala. As depicted in Figure 7, there is a positive correlation between the compile pass rate and the test pass rate, but a compile pass does not necessarily mean a test pass.
From Figure 7, it could be found that the written language of the prompts would affect the model performance. We are surprised that DeepSeek-Coder-1.5B-Instruct, proposed by a Chinese institution, is the most favorable model for English questions. On the other hand, some native English speakers, such as StarCoder2-15B-Instruct-v0.1, Open AI-o1-preview, and Open AI-o1-mini, perform better on Chinese questions.

### v3

@@ 4 Experiments
OpenCoder-1.5B-Instruct 26.05 40.03 31.50 42.64 25.17 39.12 23.75 13.97 30.16 26.67 44.12 38.30 33.52
Qwen2.5-Coder-1.5B-Instruct 18.37 34.75 29.00 33.50 28.32 41.33 17.50 15.81 40.48 23.33 47.06 28.19 30.74
DeepSeek-Coder-1.3B-Instruct 16.74 29.91 32.50 37.06 22.73 35.54 18.75 9.19 27.78 25.00 36.76 30.32 27.65
@@
Yi-Coder-9B-Chat 39.07 46.04 39.50 64.97 46.50 49.66 42.50 34.93 48.41 41.67 58.82 49.47 47.13
OpenCoder-8B-Instruct 39.53 49.12 38.00 55.58 36.01 45.92 27.50 26.47 47.62 46.67 45.59 45.74 43.63
DeepSeek-Coder-7B-Instruct-v1.5 38.37 45.16 36.00 57.36 35.66 47.96 30.00 30.88 46.03 53.33 45.59 44.15 43.48
DeepSeek-Coder-6.7B-Instruct 34.19 43.40 38.50 58.12 38.11 43.88 33.75 23.90 46.03 38.33 60.29 44.15 41.88
CodeQwen1.5-7B-Chat 36.74 44.87 46.00 51.78 29.72 40.82 26.25 24.26 42.06 41.67 48.53 44.68 40.52
CodeLlama-7B-Instruct 21.40 21.70 30.50 34.26 20.28 40.48 8.75 11.76 34.92 15.00 50.00 29.26 27.06
@@
Qwen2.5-Coder-14B-Instruct 53.26 58.50 41.00 69.54 69.23 46.26 51.25 43.01 49.21 60.00 69.12 57.45 55.28
DeepSeekCoder-v2-Lite-Instruct 45.81 57.18 38.50 56.85 52.80 44.56 42.50 33.82 52.38 33.33 50.00 51.60 48.73
StarCoder2-15B-Instruct-v0.1 38.37 42.23 29.00 59.90 37.06 40.99 42.50 28.68 54.76 33.33 42.65 45.74 41.79
CodeLlama-13B-Instruct 24.88 21.41 31.00 31.47 18.18 41.67 16.25 13.24 35.71 15.00 45.59 32.45 27.59
@@
DeepSeekCoder-v2-Instruct 52.79 63.64 43.00 71.57 75.87 47.45 46.25 52.94 53.97 51.67 63.24 59.57 58.09
Qwen2.5-Coder-32B-Instruct 51.86 60.85 43.00 73.10 69.93 47.11 55.00 44.85 56.35 61.67 61.76 60.64 56.88
DeepSeekCoder-33B-Instruct 38.37 50.59 35.50 65.99 50.00 49.49 43.75 39.71 49.21 53.33 54.41 48.40 48.61
CodeLlama-34B-Instruct 23.72 22.73 26.50 37.56 18.18 43.71 17.50 17.65 38.10 26.67 51.47 30.85 29.22
@@
Qwen2.5-72B-Instruct 52.56 61.44 43.00 66.50 76.57 48.47 55.00 51.10 52.38 51.67 55.88 55.32 56.88
Llama3.1-70B-Instruct 46.51 54.69 34.50 65.48 64.69 45.24 51.25 38.60 56.35 46.67 57.35 53.72 51.45
@@
OpenAI o1-preview 71.63 71.99 49.50 72.59 80.77 51.53 50.00 63.97 57.14 60.00 67.65 68.09 65.62
OpenAI o1-mini 70.23 75.66 41.50 71.07 81.47 48.47 56.25 59.19 54.76 60.00 69.12 67.55 64.73
Claude-35-Sonnet 61.63 65.40 53.00 71.83 77.27 48.81 53.75 63.24 58.73 68.33 64.71 68.62 62.57
GPT 4o-0806 57.21 67.60 46.00 74.37 76.92 48.47 63.75 55.88 60.32 63.33 70.59 64.89 61.77
Doubao-Coder-Preview 56.98 64.66 43.00 71.07 74.48 49.15 45.00 59.19 50.00 48.33 60.29 55.32 58.92
DeepSeek-v2.5 51.86 63.78 43.00 69.54 75.17 49.66 47.50 57.35 53.17 60.00 60.29 61.70 58.65
GLM-4-Plus 49.77 59.97 43.00 71.32 72.73 46.43 56.25 50.00 57.14 61.67 63.24 52.66 56.40
Qwen-Max 47.21 61.14 42.00 63.20 72.73 47.11 53.75 55.15 57.94 41.67 54.41 50.53 55.16
@@ 4.1 Experimental Setup
We chat with the close-sourced models via API calls similarly, where top_k is set to 1 and max_new_tokens is set to 2048. The prompt template is kept the same. After model inference, the first code block with the corresponding programming language formatted in Markdown is extracted from the generated output. If no Markdown code block is detected, a heuristic approach is employed to identify and extract incomplete code snippets. The extracted code, combined with predefined test cases, is then used to synthesize the complete code, which is evaluated for correctness.
@@ 4.2 Results and Analysis
As illustrated in Figure 5, we visualize the model performance on different domains in FullStack Bench. From Figure 5, we could find that the performance of AI coders varies significantly in BP, AP, MA, ML, SC, and MM. The largest range occurs in MA, with the best mathematician being OpenAI o1-mini (81.47) while the worst is CodeLlama-34B-Instruct (18.18). Mathematical programming requires models to be proficient in both math and code, and those trained on a code-highly-concentrated corpus would struggle to achieve high scores in MA. Similarly, the variances of ML, SC, and MM, are also remarkable, as each of these problems requires domain knowledge beyond coding. Moreover, in Figure 5, we visualize the first place in each model division. The performance trends of OpenAI-o1-preview, Qwen2.5-72B-Instruct, DeepSeek-Coder-v2-Instruct, and Qwen2.5-Coder-14B-Instruct are generally consistent. OpenAI-o1-Preview has a clear advantage on BP, AP, and SC. On ML and OS, OpenAI-o1-preview only achieves the second place. On the remaining domains, there are slight performance gaps between OpenAI-o1-preview and the runner-ups. Independent of those larger models, the performance trends of OpenCoder-1.5B-Instruct and Qwen2.5-Coder-7B-Instruct are close to each other. These two small models fall behind on BP, AP, MA, and SC, suggesting that reasoning power in these domains may only emerge in the larger models.
@@
OpenCoder-1.5B-Instruct 73.33 20.09 50.00 10.87 39.71 51.88 38.65 51.28 58.33 29.27 20.00 42.11 33.33 28.57 35.00 50.00 33.52
Qwen2.5-Coder-1.5B-Instruct 86.67 12.62 22.22 6.52 42.65 58.75 37.99 51.28 47.22 29.09 12.50 21.05 16.67 7.14 43.75 57.89 30.74
DeepSeek-Coder-1.3B-Instruct 60.00 10.28 43.94 4.35 30.88 46.88 35.15 52.56 63.89 24.35 5.00 34.21 15.00 7.14 28.75 44.74 27.65
@@
Yi-Coder-9B-Chat 76.67 25.23 75.25 13.04 41.18 67.50 42.58 74.36 55.56 45.62 27.50 63.16 51.67 46.43 56.25 65.79 47.13
OpenCoder-8B-Instruct 66.67 29.91 61.62 30.43 48.53 61.25 41.27 70.51 58.33 40.05 25.00 39.47 50.00 51.79 61.25 60.53 43.63
DeepSeek-Coder-7B-Instruct-v1.5 50.00 24.77 68.18 28.26 54.41 63.13 40.83 66.67 44.44 41.29 21.25 52.63 43.33 39.29 55.00 50.00 43.48
DeepSeek-Coder-6.7B-Instruct 76.67 23.36 56.06 13.04 52.94 65.00 41.48 65.38 50.00 39.34 31.25 47.37 43.33 28.57 57.50 60.53 41.88
CodeQwen1.5-7B-Chat 73.33 30.37 41.92 15.22 48.53 57.50 44.76 60.26 44.44 37.44 15.00 52.63 58.33 46.43 52.50 60.53 40.52
CodeLlama-7B-Instruct 76.67 13.08 43.94 11.96 23.53 54.38 32.97 37.18 27.78 21.98 17.50 21.05 20.00 25.00 41.25 50.00 27.06
@@
Qwen2.5-Coder-14B-Instruct 93.33 39.25 65.66 41.30 63.24 66.25 43.45 80.77 77.78 56.22 32.50 71.05 73.33 42.86 62.50 68.42 55.28
DeepSeekCoder-v2-Lite-Instruct 50.00 36.45 52.53 27.17 61.76 64.38 41.27 73.08 58.33 49.35 38.75 55.26 53.33 42.86 57.50 60.53 48.73
StarCoder2-15B-Instruct-v0.1 56.67 21.03 60.61 29.35 47.06 49.38 31.44 70.51 44.44 42.36 28.75 71.05 35.00 32.14 63.75 52.63 41.79
CodeLlama-13B-Instruct 63.33 11.68 50.00 15.22 33.82 55.63 33.84 35.90 36.11 21.33 17.50 42.11 25.00 21.43 38.75 47.37 27.59
@@
DeepSeekCoder-v2-Instruct 83.33 43.46 72.73 28.26 66.18 69.38 42.36 80.77 61.11 60.43 41.25 65.79 68.33 69.64 66.25 68.42 58.09
Qwen2.5-Coder-32B-Instruct 83.33 36.92 76.77 46.74 54.41 71.25 40.39 79.49 58.33 59.06 35.00 63.16 76.67 50.00 70.00 57.89 56.88
DeepSeekCoder-33B-Instruct 60.00 26.64 68.18 19.57 57.35 71.25 44.10 66.67 50.00 48.10 32.50 60.53 50.00 46.43 60.00 57.89 48.61
CodeLlama-34B-Instruct 76.67 12.62 47.98 11.96 35.29 63.75 33.62 37.18 44.44 23.34 18.75 36.84 21.67 21.43 48.75 47.37 29.22
@@
Qwen2.5-72B-Instruct 80.00 32.24 57.07 33.70 63.24 73.75 41.92 79.49 77.78 61.02 37.50 63.16 61.67 66.07 68.75 68.42 56.88
Llama3.1-70B-Instruct 76.67 28.50 72.73 40.22 50.00 77.50 28.38 71.79 41.67 54.50 36.25 65.79 58.33 46.43 68.75 57.89 51.45
@@
OpenAI o1-preview 90.00 51.40 80.81 55.43 64.71 75.63 42.36 78.21 75.00 68.19 62.50 84.21 86.67 83.93 71.25 78.95 65.62
OpenAI o1-mini 90.00 66.36 76.77 53.26 67.65 75.00 34.06 80.77 88.89 67.71 58.75 78.95 83.33 82.14 68.75 68.42 64.73
Claude-35-Sonnet 86.67 41.12 75.76 64.13 50.00 81.88 40.39 70.51 47.22 66.59 46.25 84.21 80.00 83.93 71.25 55.26 62.57
GPT 4o-0806 93.33 50.47 72.73 28.26 67.65 71.25 43.23 83.33 44.44 64.81 52.50 73.68 85.00 73.21 76.25 57.89 61.77
Doubao-Coder-Preview 76.67 40.65 74.24 43.48 52.94 66.25 37.77 84.62 77.78 62.74 45.00 71.05 66.67 75.00 66.25 65.79 58.92
DeepSeek-v2.5 86.67 45.33 74.75 30.43 61.76 76.25 40.83 75.64 91.67 60.90 35.00 68.42 65.00 64.29 66.25 71.05 58.65
GLM-4-Plus 83.33 38.79 74.24 32.61 54.41 75.63 41.70 80.77 80.56 58.29 31.25 52.63 60.00 58.93 72.50 55.26 56.40
Qwen-Max 76.67 32.71 40.40 28.26 61.76 73.75 42.79 85.90 52.78 60.37 35.00 60.53 76.67 46.43 72.50 52.63 55.16
@@ 4.3 Analysis on the performance of different programming languages
We present results of different programming languages in Table 3 and Figure 6. Except for Java, the performance of AI coders on different languages has significant variance. The performance gaps are relatively larger in C#, D, PHP, Ruby, Rust, and Scala. Besides, as our SandboxFusion provides feedback from the compilers, we evaluate the compilation pass rates of model responses in compiled languages including C++, D, Java, Rust, and Scala. In Figure 7, there is a positive correlation between the compile pass rate and the test pass rate, but a compile pass does not necessarily mean a test pass. Moreover, from Figure 7, we found that the written language of the prompts would affect the model performance. We are surprised that DeepSeek-Coder-1.5B-Instruct, proposed by a Chinese institution, is the most favorable model for English questions. On the other hand, some native English speakers, such as StarCoder2-15B-Instruct-v0.1, Open AI-o1-preview, and Open AI-o1-mini, perform better on Chinese questions.